OpenStudy (anonymous):

solve 6=x+2/3. show your steps.

OpenStudy (anonymous):

6-2/3=x 5 3/3 -2/3=x 5 1/3=x i guess thats the answer

OpenStudy (gorv):

6-2/3=x 6-0.666=x 5.333=x

OpenStudy (anonymous):

medals are accepted as well ;)

OpenStudy (doc.brown):

\[x+\frac{2}{3}=6\]\[x+\frac{2}{3}-\frac{2}{3}=6-\frac{2}{3}\]\[x=6-\frac{2}{3}\]\[x=\frac{18}{3}-\frac{2}{3}\]\[x=\frac{16}{3}\]